OpenStudy (anonymous):

Solve 5^(x + 5) = 9^x. Please Explain how the answer was determined! Thank you! :)

OpenStudy (anonymous):

\[x^{a+b} = x^a * x^b\] we have:\[5^x * 5^5 = 9^x\]\[5^5 = \left( \frac{ 9 }{ 5 } \right)^x = 3125\] can you solve for x now?

OpenStudy (anonymous):

I'm having trouble with solving exponential equations with unequal bases. My best bet at solving this problem would be with your help (if you want to provide it). :) So to answer your question I cannot solve for x without your help. :)

OpenStudy (tkhunny):

Have you met the logarithm?

OpenStudy (anonymous):

\[x = \frac{ \log(3125) }{ \log(9/5) }\]